Find the equation, in standard form, with all integer coefficients, of the line perpendicular to 3x-4y=9 and passing through (-2,-1)
Answer by zeynep(43)

the formula for the slope of a line is; -(coefficient of x/coefficient of y)
so the slope of the line, 3x-4y=9, is ; 3/4.
the product of the slopes of two perpendicular lines is -1.
Therefore, the slope of the new line is -1/(3/4)=-4/3
A line going through the point (a,b) and having slope of m would have the equation;
y-a=m(x-b)
Therefore the equation for our new line is;
y+1=-4/3(x+2)
y+1=-4/3x-8/3 (if we multiply both sides by 3)
3y+3=-4x-8
4x+3y+11=0